OpenStudy (anonymous):

I learned this earlier in the year and never was good at it, but i have to retake this assignment. Here is he first problem (: I am expected to show my work. The sides of a triangle have measures of x, x, and y, and the perimeter of the triangle is 46. The measure y is 4 centimeters longer than the measure x. Write two equations which represent the two sentences. Solve the system you wrote for this using any technique.

OpenStudy (anonymous):

x+x+y = 46 y=4+x

OpenStudy (anonymous):

plugging y = 4+x in eqn 1, x+x+4+x = 46 3x=42 x=14 hence y=4+x = 4+14=18

OpenStudy (anonymous):

Thanks buddyyyy (:

OpenStudy (anonymous):

Why is this stated as "the measure y is 4 cm longer than the measure x?" Why not "y is 4 cm longer than x?"

OpenStudy (anonymous):

I have no idea i only copied and pasted ^_^